You have been asked to design a high performing and highly redundant storage array with a minimum of 32 TB of usable space for files. 4 TB hard drives cost $200, 6 TB hard drives cost $250, 8 TB hard drives cost $300, and 10 TB hard drives cost $350. Explain which type of RAID you would choose and the quantity and types of drives you would use for your solution.
Answer #1

Number of 4 TB hard drive need = 32/4 = 8 and cost = 200*8 = 1600 $

Number of 6 TB hard drive need = max(32/6) = 6 and cost = 250*6 = 1500 $

Number of 8 TB hard drive need = 32/8 = 4 and cost = 300*4 = 1200 $

Number of 10 TB hard drive need = 32/10 = 4 and cost = 350*4 = 1400 $

Hence using 4 hard drive of 8 TB will minimize the cost